### v1

@@ 2.5 A note on analytic solutions
The analytic solutions are applicable under the assumption of a homogeneous viscosity mechanism. These models can serve as benchmarks for testing the behavior of numerical schemes. However, in real astrophysical discs, a region where the viscosity can be considered homogeneous is in most cases not only limited, but also changes in size on a time scale comparable to the viscous time scale. In accretion discs surrounding stellar-mass compact objects, viscosity due to MRI-driven turbulence is believed to be suppressed in cooler recombined regions away from the gravitational center. Cold, massive discs — such as protoplanetary discs or found in active galactic nuclei — are often characterized as multizone structures, where the viscosity due to gravitoturbulence depends on the details of thermal balance and external irradiation.
@@ 3 Angular Momentum Transport by Magnetic Fields
Accretion can only happen in a rotationally-supported disc if there is some mechanism for extracting angular momentum from the rotating fluid elements. This is why the Shakura-Sunyaev ansatz (1) has been so essential in building analytic models of accretion discs. Over the last fifty years, significant progress has been made in understanding the basic physics of this angular momentum-transporting stress, and the advent of large-scale numerical simulations has created the opportunity for building models that incorporate this basic physics.
While nonaxisymmetric spiral waves can contribute to hydrodynamically transporting angular momentum outwards (see, e.g., section 5 of [76] for a detailed review), it appears that in discs with high microscopic electrical conductivity and high orbital Mach numbers, magnetic stresses will dominate [77, 78]. There are three mechanisms whereby magnetic fields can transport angular momentum, which are not mutually exclusive: magnetorotational (MRI) turbulence, mean-field magnetic stresses within the disc, and magnetocentrifugal winds. [^8] There is a semantic issue here in that it can be argued that any form of magnetic angular momentum transport is magnetorotational in nature. We are here drawing a distinction between local turbulence and more global mean field transport. It is only the former that might be describable in terms of the local Shakura-Sunyaev prescription (1).
@@ 3.1 Magnetorotational (MRI) Turbulence
A major breakthrough in accretion disc theory occurred in the early 1990’s with the realization that the magnetorotational instability (MRI) [33, 79, 80, 81] provides a robust mechanism for exciting turbulence in a disc with precisely the necessary properties to both transport angular momentum outward between fluid elements and to dissipate mechanical energy into internal energy [82]. An excellent review of the physics of this instability is Balbus and Hawley 83. The simplest version consists of axisymmetric perturbations on a vertical magnetic field, a situation that is always unstable in the presence of negative radial angular velocity gradients provided that the field is weak enough that the vertical Alfvén wave period is longer than of order the orbital period. Because the wave period is proportional to the vertical wavelength, we require that the minimum unstable mode wavelength $\sim v_{{\rm A}z}/\Omega$ can fit within some measure of the size of the disc, particularly the disc scale height. Here $v_{{\rm A}z}$ is the local vertical Alfvén speed and $\Omega$ is the local angular velocity. If the disc is supported vertically by thermal pressure, then this requires the magnetic pressure due to the vertical magnetic field to be subthermal, i.e. plasma betas greater than unity. The growth rate can be greatly reduced by curvature of a strong toroidal field when the toroidal Alvén speed exceeds of order the geometric mean of the sound and orbital speeds [84, 85]. However, even a purely toroidal magnetic field will be unstable to transient growth of nonaxisymmetric perturbations in the presence of negative radial angular velocity gradients [81], and here stability only occurs if the Alfvén speed is less than the orbital speed [86, 87]. [^9] Exponential growth of true nonaxisymmetric unstable modes can only exist in a rotating shear flow when one solves the global eigenvalue problem with boundary conditions. Unfortunately, linear analyses (axisymmetric or nonaxisymmetric) beyond the simplest local instability can be complicated, particularly when they account for the vertical or global structure of the disc, and/or the coupling to other instabilities [88, 86, 87, 89, 85, 90].
@@
Over the years, substantial effort by many groups has been made to understand the nonlinear outcome of the MRI in three computational domains: local shearing box simulations of a rotating patch of the disc [91] that do not include (vertically unstratified) or do include (vertically stratified) vertical gravity, and global simulations. These simulations start with a laminar flow with some initial magnetic field that then becomes unstable and turbulent. The vertically stratified and global simulations of weak field MRI exhibit dynamo behavior in which the radial and toroidal fields alternate polarity and rise with height in the disc [92] - the so-called MRI butterfly diagram. An excellent modern review of the dynamo properties of MRI turbulence can be found in section 5.3 of [93]. Nevertheless, the topology of the initial magnetic field tends to have a long term effect on the properties of the turbulence and the turbulent stresses. For example, in shearing box simulations, the saturated turbulent stress increases above a certain level with increasing net vertical magnetic flux until the MRI is stabilized [91, 94]. In global simulations the long term outcome tends to depend on how much local vertical flux is present. This dependence on initial field topology is unfortunate given that we have little information concerning the magnetic field geometry in the material that fuels accretion discs across astrophysics. In recent work, however, vertical flux was self-consistently generated by a dynamo operating on an initially purely toroidal field configuration [95].
@@
The resistive and viscous dissipation scales in the turbulent cascade are often impossible to resolve with existing computational resources, and even getting enough dynamic range to produce a convincing inertial range has only recently been achieved [96]. Hence the vast majority of simulations are run in such a manner as to have magnetic and kinetic energy dissipate at the grid scale - so-called ILES (Implicit Large Eddy Simuations) methods. Unfortunately, this is also not without consequences. Zero net flux, vertically unstratified shearing box ILES simulations with simplified (isothermal) thermodynamics of MRI turbulence do not converge: the saturation level of the turbulent stresses decreases with increasing numerical resolution [94, 97]. This also appears to be the case when vertical stratification is included [98]. However, when explicit viscosity ($\nu$) and resistivity ($\eta$) (with values much larger than are realistic) are included, then numerical convergence is obtained with zero net magnetic flux [99, 100]. The resulting stresses exhibit a strong dependence on magnetic Prandtl number $\mathrm{Pm}\equiv\nu/\eta$. This dependence can even result in thermal/viscous instabilities with stable upper and lower temperature branches. This has been explored in the context of the inner regions of black hole X-ray binary discs by [101]. A compelling explanation for the Prandtl number dependence in zero net flux, unstratified shearing boxes has been recently presented in [102]. This has also been extended to vertically stratified shearing boxes at high magnetic Prandtl number (the regime of relevance to luminous active galactic nuclei and the inner regions of X-ray binaries) in [103, 104].
@@
Provided explicit dissipation coefficients are included and the box size is sufficiently large, unstratified zero net flux simulations even produce saturated turbulent stresses that are proportional to the thermal pressure [105], in agreement with the alpha prescription. This dependence is weakened, however, in the presence of net magnetic flux through the box.
@@
Thermodynamics can also affect the properties of MRI turbulence. In a radiation pressure dominated environment, such as in the innermost region of a standard Shakura-Sunyaev disc, photon diffusion can cause the turbulence to be highly compressible even though the turbulent fluid speeds are much less than the total sound speed in the radiation-dominated plasma [106]. This is because photon diffusion out of a compressing region reduces the build-up of radiation pressure to resist the compression. This diffusion is also accompanied by significant damping of compressible motions [107], i.e. a large radiation bulk viscosity. As a result, the turbulence has a high effective magnetic Prandtl number, and the ratio of Maxwell to turbulent Reynolds stress is enhanced [108].
@@
In addition to mediating the dissipation of accretion power through a turbulent cascade, MRI turbulence can also affect the cooling rate of an accretion disc in the radiation pressure dominated regime. The large density variations in the compressible turbulence create porosity that enhances the photon diffusion rate, with photons diffusing mostly through low density channels. (This effect has been quantified in compressible convective turbulence in massive stars by [109].) In addition, buoyant magnetic field concentrations that form in the turbulence can produce bulk transport of photons through vertical advection [110, 111].
@@
Vertically stratified shearing box simulations of MRI turbulence that incorporate radiation transfer with temperature-dependent opacities can become vertically unstable to convection when the opacities are high enough [112, 113, 114, 115]. Depending on the behavior of the opacity with temperature, the simulations can alternate between convective and radiative cycles or maintain persistent convection. When convection is present, the alternating polarity of azimuthal field so characteristic of the MRI turbulent dynamo in vertically stratified shearing boxes [92] is suppressed, likely by vertical mixing of field [114]. In some (but not all [115]) cases, this convection can enhance the time-averaged MRI turbulent stresses even in the absence of net vertical magnetic flux. All of this behavior has also been observed in global radiation MHD simulations of MRI turbulence with temperature-dependent opacities [116]. The enhancement of stresses appears to be associated with the creation of vertical magnetic field by convection [112], but more research would be helpful to fully understand this. Recent numerical experiments using non-radiative shearing box simulations but with fixed resistivity and fixed optically thin cooling time in the uppermost layers (to control the vertical entropy gradient) find that the flow can alternate between convection and MRI-dominated episodes, with an enhancement of stress during the MRI phases (with little or no convection at those times) compared to simulations in which convection never takes place [117].
@@
Although MRI is generally thought to be a weak field instability, simulations with substantial vertical flux can lead to configurations with strong turbulence and in which toroidal magnetic pressure dominates thermal pressure and provides hydrostatic support of the disc against vertical gravity [118, 119, 120]. The regular toroidal field reversals so characteristic of the weak field MRI dynamo (the butterfly diagram) are lost in this regime. The vertical structure of such discs appears to be consistent with some of the analytic ideas proposed for magnetically dominated discs [12, 13, 15], and it is likely that such discs are thermally and viscously stable.
@@
Strongly magnetized, turbulent structures can also arise in simulations that start with no vertical magnetic field, provided a sufficiently strong toroidal field is present. This was perhaps first anticipated by [121], who proposed a dynamo mechanism coupling the Parker instability, MRI, and magnetic reconnection. That such a thing is possible was first demonstrated numerically by [122]: a strong toroidal field that vertically supports the disc is Parker unstable, creating vertical magnetic field that is MRI unstable. The nonaxisymmetric MRI on the toroidal field (which as noted above does not require a weak field) also plays a role in the resulting nonlinear state. Conducting boundary conditions that preclude vertical field escape were used for these simulations [122], and simulations with outflowing boundary conditions, both local [120] and global [123] were unable to retain strong toroidal fields and replicate this result. On the other hand, various forms of outflow boundary conditions were implemented in recent vertically stratified shearing box simulations with pure toroidal fields [124], and these succeeded in producing self-sustaining strongly magnetized turbulent configurations, like those in the original work of [122]. Without vertical field, the midplane must be magnetically dominated in the saturated state for this to happen, otherwise buoyant field loss returns the system to a weakly magnetized state of MRI turbulence. Much more work needs to be done to understand the origin of this strongly magnetized regime and how it might play out in different astrophysical environments.
@@ 3.2 Mean-Field Magnetic Stresses Within the Disc
Depending on initial conditions, global simulations of magnetized accretion flows can lead to states where magnetic pressure dominates thermal pressure. These tend to have large scale, coherent magnetic fields [125, 126, 127], in which mean Maxwell stresses are comparable to or even dominate the outward angular momentum transport compared to turbulent Maxwell stresses, i.e. $<B_{r}B_{\phi}>\sim<B_{r}><B_{\phi}>$. Perhaps the most extreme version of this regime has come from recent zoom-in cosmological simulations of the fueling of a supermassive black hole [128]. Midplane plasma beta’s between $10^{-6}$ and $10^{-2}$ exist in the disc on a scale of $\sim 100$ gravitational radii. The field is predominately toroidal and is maintained by inward advection of magnetic flux. The Alfvén speed greatly exceeds the geometric mean of the sound and orbital speeds, which as we discussed above is the limit where MRI growth rates are significantly reduced [84, 85]. Nevertheless, strong turbulence exists in the disc, which is supported vertically by both magnetic pressure and turbulent kinetic energy gradients [127]. Analytic scalings of these simulation results can be found in [129].
A hybrid situation has also been found in simulations with net poloidal flux in which most of the accretion occurs in a magnetically dominated “corona” that sandwiches an equatorial, thermal pressure-dominated, MRI turbulent disc, and where angular momentum transport in the corona is dominated by mean field Maxwell stresses [130, 131, 132]. This high altitude flow can even consist of alternating layers of turbulent and mean field Maxwell stresses, with the turbulent layers being due to MRI [133].

### v2

@@ 2.5 A note on analytic solutions
The analytic solutions are applicable under the assumption that viscosity is a power law function of radius and surface density. These models can serve as benchmarks for testing the behavior of numerical schemes. However, in real astrophysical discs, a region where the viscosity can be considered to be smoothly varying is in most cases not only limited, but also changes in size on a time scale comparable to the viscous time scale. In accretion discs surrounding stellar-mass compact objects, viscosity due to MRI-driven turbulence is believed to be suppressed in cooler recombined regions away from the gravitational center. Cold, massive discs — such as protoplanetary discs or those found in active galactic nuclei — are often characterized as multizone structures, where the viscosity due to gravitoturbulence depends on the details of thermal balance and external irradiation.
@@ 3 Angular Momentum Transport by Magnetic Fields
Accretion can only happen in a rotationally-supported disc if there is some mechanism for extracting angular momentum from the rotating fluid elements. This is why the Shakura-Sunyaev ansatz (1) has been so widely used in building analytic models of accretion discs. Over the last fifty years, significant progress has been made in understanding the basic physics of this angular momentum-transporting stress, and the advent of large-scale numerical simulations has created the opportunity for building models that incorporate this basic physics.
While nonaxisymmetric spiral waves can contribute to hydrodynamically transporting angular momentum outwards (see, e.g., section 5 of pap95 for a detailed review), it appears that in discs with high microscopic electrical conductivity and high orbital Mach numbers, magnetic stresses will dominate [ju16, ju17]. There are three mechanisms whereby magnetic fields can transport angular momentum, which are not mutually exclusive: magnetorotational (MRI) turbulence, mean-field magnetic stresses within the disc, and magnetocentrifugal winds. [^9] There is a semantic issue here in that it can be argued that any form of magnetic angular momentum transport is magnetorotational in nature. We are here drawing a distinction between local turbulence and more global mean field transport. It is only the former that might be describable in terms of the local Shakura-Sunyaev prescription (1).
@@ 3.1 Magnetorotational (MRI) Turbulence
A major breakthrough in accretion disc theory occurred in the early 1990’s with the realization that the magnetorotational instability [MRI; bal91, haw91, haw92, bal92] provides a robust mechanism for exciting turbulence in a disc with precisely the necessary properties to both transport angular momentum outward between fluid elements and to dissipate mechanical energy into internal energy [bal99]. An excellent review of the physics of this instability can be found in bal98. The simplest version consists of axisymmetric perturbations on a vertical magnetic field, a situation that is always unstable in the presence of negative radial angular velocity gradients provided that the field is weak enough that the vertical Alfvén wave period is longer than of order the orbital period. Because the wave period is proportional to the vertical wavelength, we require that the minimum unstable mode wavelength $\sim v_{{\rm A}z}/\Omega$ can fit within some measure of the size of the disc, particularly the disc scale height. Here $v_{{\rm A}z}$ is the local vertical Alfvén speed and $\Omega$ is the local angular velocity. If the disc is supported vertically by thermal pressure, then this requires the magnetic pressure due to the vertical magnetic field to be subthermal, i.e. plasma betas greater than unity. The growth rate can be greatly reduced by curvature of a strong toroidal field when the toroidal Alvén speed exceeds of order the geometric mean of the sound and orbital speeds [pes05, das18]. However, even a purely toroidal magnetic field will be unstable to transient growth of nonaxisymmetric perturbations in the presence of negative radial angular velocity gradients [bal92, fog95, ter96]. [^10] Exponential growth of true nonaxisymmetric unstable modes can only exist in a rotating shear flow when one solves the global eigenvalue problem with boundary conditions. Unfortunately, linear analyses (axisymmetric or nonaxisymmetric) beyond the simplest local instability can be complicated, particularly when they account for the vertical or global structure of the disc, and/or the coupling to other instabilities [gam94, fog95, ter96, cur96, das18, kim00].
Figure 4: Snapshot of the three dimensional structure of mass density (top) and radiation energy density (bottom) of a global simulation of MRI turbulence in a disk around a supemassive black hole [jia20]. In addition to MHD, this simulation incorporates angle-dependent radiation transport with frequency-averaged (grey) opacities.
@@
Over the years, substantial effort by many groups has been made to understand the nonlinear outcome of the MRI in three computational domains: local shearing box simulations of a rotating patch of the disc [haw95] that do not include (vertically unstratified) or do include (vertically stratified) vertical gravity, and global simulations (see Fig. 4 for an example). These simulations start with a laminar flow with some initial magnetic field that then becomes unstable and turbulent. The vertically stratified and global simulations of weak field MRI exhibit dynamo behavior in which the radial and toroidal fields alternate polarity and rise with height in the disc [bra95] — the so-called MRI butterfly diagram. An excellent modern review of the dynamo properties of MRI turbulence can be found in section 5.3 of rin19. Nevertheless, the topology of the initial magnetic field tends to have a long term effect on the properties of the turbulence and the turbulent stresses. For example, in shearing box simulations, the saturated turbulent stress increases above a certain level with increasing net vertical magnetic flux until the MRI is stabilized [haw95, pes07]. In global simulations the long term outcome tends to depend on how much local vertical flux is present. This dependence on initial field topology is unfortunate given that we have little information concerning the magnetic field geometry in the material that fuels accretion discs across astrophysics. In recent work, however, vertical flux was self-consistently generated by a dynamo operating on an initially purely toroidal field configuration [jac24].
@@
The resistive and viscous dissipation scales in the turbulent cascade are often impossible to resolve with existing computational resources, and even getting enough dynamic range to produce a convincing inertial range has only recently been achieved [kaw24]. Hence the vast majority of simulations are run in such a manner as to have magnetic and kinetic energy dissipate at the grid scale - so-called ILES (Implicit Large Eddy Simulations) methods. Unfortunately, this is also not without consequences. Zero net flux, vertically unstratified shearing box ILES simulations with simplified (isothermal) thermodynamics of MRI turbulence do not converge: the saturation level of the turbulent stresses decreases with increasing numerical resolution [pes07, fro07a]. This also appears to be the case when vertical stratification is included [rya17]. However, when explicit viscosity ($\nu$) and resistivity ($\eta$) (with values much larger than are realistic) are included, then numerical convergence is obtained with zero net magnetic flux [fro07b, les07]. The resulting stresses exhibit a strong dependence on magnetic Prandtl number $\mathrm{Pm}\equiv\nu/\eta$. This dependence can even result in thermal/viscous instabilities with stable upper and lower temperature branches. This has been explored in the context of the inner regions of black hole X-ray binary discs by pot17. A compelling explanation for the Prandtl number dependence in zero net flux, unstratified shearing boxes has been recently presented by mam20. This has also been extended to vertically stratified shearing boxes at high magnetic Prandtl number (the regime of relevance to luminous active galactic nuclei and the inner regions of X-ray binaries) by hel22b, hel24.
@@
Provided explicit dissipation coefficients are included and the box size is sufficiently large, unstratified zero net flux simulations even produce saturated turbulent stresses that are proportional to the thermal pressure [ros16], in agreement with the alpha prescription. This dependence is weakened, however, in the presence of net magnetic flux through the box.
@@
Thermodynamics can also affect the properties of MRI turbulence. In a radiation pressure dominated environment, such as in the innermost region of a standard Shakura-Sunyaev disc, photon diffusion can cause the turbulence to be highly compressible even though the turbulent fluid speeds are much less than the total sound speed in the radiation-dominated plasma [tur03]. This is because photon diffusion out of a compressing region reduces the build-up of radiation pressure to resist the compression. This diffusion is also accompanied by significant damping of compressible motions [ago98], i.e. a large radiation bulk viscosity. As a result, the turbulence has a high effective magnetic Prandtl number, and the ratio of Maxwell to turbulent Reynolds stress is enhanced [jia13a].
@@
In addition to mediating the dissipation of accretion power through a turbulent cascade, MRI turbulence can also affect the cooling rate of an accretion disc in the radiation pressure dominated regime. The large density variations in the compressible turbulence create porosity that enhances the photon diffusion rate, with photons diffusing mostly through low density channels. (This effect has been quantified in compressible convective turbulence in massive stars by sch20.) In addition, buoyant magnetic field concentrations that form in the turbulence can produce bulk transport of photons through vertical advection [bla11, jia14].
@@
Vertically stratified shearing box simulations of MRI turbulence that incorporate radiation transfer with temperature-dependent opacities can become vertically unstable to convection when the opacities are high enough [hir14, hir15, col17, sce18a]. Depending on the behavior of the opacity with temperature, the simulations can alternate between convective and radiative cycles or maintain persistent convection. When convection is present, the alternating polarity of azimuthal field so characteristic of the MRI turbulent dynamo in vertically stratified shearing boxes [bra95] is suppressed, likely by vertical mixing of field [col17]. In some [but not all, sce18a] cases, this convection can enhance the time-averaged MRI turbulent stresses even in the absence of net vertical magnetic flux. All of this behavior has also been observed in global radiation MHD simulations of MRI turbulence with temperature-dependent opacities [jia20]. The enhancement of stresses appears to be associated with the creation of vertical magnetic field by convection [hir14], but more research would be helpful to fully understand this. Recent numerical experiments using non-radiative shearing box simulations but with fixed resistivity and fixed optically thin cooling time in the uppermost layers (to control the vertical entropy gradient) find that the flow can alternate between convection and MRI-dominated episodes, with an enhancement of stress during the MRI phases (with little or no convection at those times) compared to simulations in which convection never takes place [hel21]. Note that in all these cases, the angular momentum transport appears to still be fundamentally magnetic in nature. Non-magnetic hydrodynamic simulations of convection in disks tend to give rise to much smaller transport, with $\alpha$ ranging from $10^{-6}$ to $10^{-5}$ [hel18]. Moreover, early simulations actually found the transport to be inward, not outward [cab96, sto96]. Only at sufficiently high Rayleigh numbers does direct angular momentum transport by hydrodynamic convection become outward [les10, hel18].
@@
Although MRI is generally thought to be a weak field instability, simulations with substantial vertical flux can lead to configurations with strong turbulence and in which toroidal magnetic pressure dominates thermal pressure and provides hydrostatic support of the disc against vertical gravity [bai13, sal16a, sal16b]. Whether this is still the same as standard MRI turbulence is unclear: the regular toroidal field reversals so characteristic of the weak field MRI dynamo (the butterfly diagram) are lost in this regime. The vertical structure of such discs appears to be consistent with some of the analytic ideas proposed for magnetically dominated discs [par03, beg07, beg15], and it is likely that such discs are thermally and viscously stable.
@@
Strongly magnetized, turbulent structures can also arise in simulations that start with no vertical magnetic field, provided a sufficiently strong toroidal field is present. This was perhaps first anticipated by tou92, who proposed a dynamo mechanism coupling the Parker instability, MRI, and magnetic reconnection. That such a thing is possible was first demonstrated numerically by joh08: a strong toroidal field that vertically supports the disc is Parker unstable, creating vertical magnetic field that is MRI unstable. The nonaxisymmetric MRI on the toroidal field (which does not require a weak field) also plays a role in the resulting nonlinear state. Conducting boundary conditions that preclude vertical field escape were used for these simulations [joh08], and simulations with outflowing boundary conditions, both local [sal16b] and global [fra17] were unable to retain strong toroidal fields and replicate this result. On the other hand, various forms of outflow boundary conditions were implemented in recent vertically stratified shearing box simulations with pure toroidal fields [squ24], and these succeeded in producing self-sustaining strongly magnetized turbulent configurations, like those in the original work of joh08. Without vertical field, the midplane must be magnetically dominated in the saturated state for this to happen, otherwise buoyant field loss returns the system to a weakly magnetized state of MRI turbulence. Much more work needs to be done to understand the origin of this strongly magnetized regime and how it might play out in different astrophysical environments.
@@ 3.2 Mean-Field Magnetic Stresses Within the Disc
Depending on initial conditions, global simulations of magnetized accretion flows can lead to states where magnetic pressure dominates thermal pressure. These tend to have large scale, coherent magnetic fields [mac06, gab12, hop24b], in which mean Maxwell stresses are comparable to or even dominate the outward angular momentum transport compared to turbulent Maxwell stresses, i.e. $\langle B_{r}B_{\phi}\rangle\sim\langle B_{r}\rangle\langle B_{\phi}\rangle$. Perhaps the most extreme version of this regime has come from recent zoom-in cosmological simulations of the fueling of a supermassive black hole [hop24a]. Midplane plasma beta’s between $10^{-6}$ and $10^{-2}$ exist in the disc on a scale of $\sim 100$ gravitational radii. The field is predominately toroidal and is maintained by inward advection of magnetic flux. The Alfvén speed greatly exceeds the geometric mean of the sound and orbital speeds, which as we discussed above is the limit where MRI growth rates are significantly reduced [pes05, das18]. Nevertheless, strong turbulence exists in the disc, which is supported vertically by both magnetic pressure and turbulent kinetic energy gradients [hop24b]. Analytic scalings of these simulation results can be found in hop24c.
A hybrid situation has also been found in simulations with net poloidal flux in which most of the accretion occurs in a magnetically dominated “corona” that sandwiches an equatorial, thermal pressure-dominated, MRI turbulent disc, and where angular momentum transport in the corona is dominated by mean field Maxwell stresses [suz14, zhu18, mis20]. This high altitude flow can even consist of alternating layers of turbulent and mean field Maxwell stresses, with the turbulent layers being due to MRI [jac21].